2. An electrical company claims that the average life of the bulbs it manufactures is 1 200 hours with a standard deviation of 250 hours. If a random sample of 100 bulbs is chosen, what is the probability that the sample mean will be less than 1150 hours? (5 points)


a. Z = - 2

b. Z = - 3

c. Z = - 4

d. Z = - 5


1
Expert's answer
2022-04-27T16:24:32-0400

Let "X=" the sample mean: "X\sim N(\mu,\sigma^2/n)"


"P(X<1150)=P(Z<\dfrac{1150-1200}{250/\sqrt{100}})"

"=P(Z<-2)\approx0.022750"